The doctor orders ʒ iiss of Maalox po 30 minutes ac and hs. The stock supply is an 8 oz bottle of Maalox. How many ounces will the patient receive in 24 hours?
Answer #1

Doctors order 3 iiss

3 resembles the symbol of scruples

ii - 2

ss- 1/2

3 iiss = 21/2 scruples

1 scruple = 20 grains

2 1/2 scruple = 20 X 2 .5

2 1/2 scruple = 50 grains

50 grains = 0.11 ounce (divide the value by 437.5)

Doctors ordered ac and hs (twice in a day)

0.11 X 2 = 0.22 ounces per 24 hour.

Hence the patient receives 0.22 ounces in 24 hours.